OpenStudy (anonymous):

Is this statement true or false and why: f(x) = 1^x is not an exponential function

OpenStudy (anonymous):

no because it is identically 1

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

i think it is an exponential function which is defined as a function b^x where b is any real number

OpenStudy (anonymous):

no exponential function of the form \[f(x)=a\times b^x\] restrictions are \[b>0\] and \[b\neq 1\]

OpenStudy (anonymous):

\[f(x)=1^x\] same as \[f(x)=1\] a constant function, not exponential

OpenStudy (anonymous):

so is it true or false?

OpenStudy (anonymous):

true because it says "NOT and exponential function'

OpenStudy (anonymous):

wait so why is it true?